Recent Log Entries

COMMENT ON BO dipole magnet mapping - SAD 2023

Log Number: 4159087Logbook(s): Body: 

Magnet Measurement:

We have completed all BO dipole magnetic field mapping and are putting together mapping data now, that will be ready next week and will be sent to Matt when completed.

Possession of magnet is now in Matt/Injector group hands.

COMMENT ON ZA dipoles return LCW water line replacement (7 magnets) - SAD 2023

Log Number: 4159086Logbook(s): Body: 

Magnet Measurement:

We have gathered all old hoses from around machine from ZA dipole hose replacement. All tools and related items picked up. Magnet work areas cleaned up. All hoses are staged in South downstairs pit area for RAD check, after approval we will bring hoses to Magnet Shop.

We are gathering hose lengths for a few locations that need a replacement 1/2'' LCW return main feed line. That work will continue next week.

2L23 leak situation

Log Number: 4159085Logbook(s): Tags: Entry Makers: beneschBody: 

Foreline pressure spiked high enough that I had an operator turn off the turbo. I then went to talk to cryo. After discussion, Joe Wilson decided to turn the liquid level set point down to 25% for the weekend. Figure 2 has turbo pump power, liquid level and pump temperature. Lower pane has foreline pressure and a Penning ? gauge from the SRF SL vacuum page. One sees on that gauge that the pressure was below 1E-4 Torr before the cooldown started, so superinsulation would have worked. In the upper pane one sees that the Agilent pumps were on during the full episode.

2L23 LL set to 25%

Log Number: 4159083Logbook(s): Entry Makers: radovicBody: 

Set 2L23 LL to 25% in response to helium to insulating vacuum leak, after a conversation with Jay Benesch and Joe Wilson. Changed low LL alarm set point, as well. The plan is to leave it like that over the weekend and then re-evaluate the situation on Tuesday.

COMMENT ON Flowswitch to flowmeter conversion

Log Number: 4159081Logbook(s): Body: 

Wired jumpers on cross connect and installed assembly in rack IN01B03. Routed ADC cable through racks and staged next to VME card on iocnl6 in IN01B05. Still need to throw some permanent labels on cables. Ready for testing / hot checkout.

A3 PSS hardware seems good, machine shop modifying the hardware on A4

Log Number: 4159078Logbook(s): Body: 

Jerry and I think A3 will function properly as a safety device, with IN and OUT positions registering properly, as required. On Tuesday, we can add air pressure and make sure all compression fittings seal properly, pressure gauges read expected values.

A4 hardware needs a few modifications at the machine shop. Mods will ensure IN/OUT positions register properly.

RF HEC Restart Checklist

Log Number: 4159077Logbook(s): Entry Makers: gelhaarwissmannBody: 

Completed the HEC Restart Checklists for C20/50/75 HPA/HVPS, C100 HPA/HVPS, and Separator IOT/HVPS.

chopper water skid is ON and at temp

Log Number: 4159076Logbook(s): Body: 

Great work my Mark and Larry and Leo, the chopper water skid is ON again. We can see it on the ion pumps, with current increase from 5 to 15 uA in chopper cavities, and 1 to 2uA near buncher. On Tuesday, we will start the process of exercising the the cavities with power at levels required for 200keV beam. The pump cart can eat gas load if it's substantial.

iocitfrffcc2 restoration failed

Log Number: 4159067Logbook(s): Tags: Body: restorelib 2.6 Target restore date was 2023-04-29 07:13:51 The output from the myRestore execution was as follows: iocitfrffcc2 Rf12Fcc_itf Pause: 60 second(s) Execute: /usr/opsuser/opshome/burt/REBOOT/Rf12Fcc/DLOAD_ICAL.csh iocitfrffcc2 (success) c100_rffcc_itf@2023-04-29 07:13 35(35) restored c100_ecal@2023-04-29 07:13 19(19) restored Execute: /usr/opsuser/opshome/burt/REBOOT/Rf12Fcc/set_IFPWP_timestamp.csh M12 (success) Execute: myRestore -execute -P c100_ecal -t RM12_ECAL -m BLDGNUM=M ZONE=1 CAV=2 (success) c100_ecal@2022-07

iocitfrffcc1 restoration failed

Log Number: 4159066Logbook(s): Tags: Body: restorelib 2.6 Target restore date was 2023-04-29 07:13:49 The output from the myRestore execution was as follows: iocitfrffcc1 Rf12Fcc_itf Pause: 60 second(s) Execute: /usr/opsuser/opshome/burt/REBOOT/Rf12Fcc/DLOAD_ICAL.csh iocitfrffcc1 (success) c100_rffcc_itf@2023-04-29 07:13 35(35) restored c100_ecal@2023-04-29 07:13 19(19) restored Execute: /usr/opsuser/opshome/burt/REBOOT/Rf12Fcc/set_IFPWP_timestamp.csh M11 (success) Execute: myRestore -execute -P c100_ecal -t RM11_ECAL -m BLDGNUM=M ZONE=1 CAV=1 (success) c100_ecal@2022-07

MBO0I06 dipole magnet for 1D keV spectrometer line now installed at injector

Log Number: 4159063Logbook(s): Body: 

The MBO0I06 dipole magnet for 1D keV spectrometer line is now on the beamline and ready for alignment. Mike B. and Joe M. will provide field map and the comparison to the old map before shims were added to the return legs. This will help us estimate current required to bend 200 keV beam. S&A can align the 1D line when they are ready.

girder3 phase2 beamline leak check OK

Log Number: 4159062Logbook(s): Body: 

The ion pump on the 1D spectrometer beamline (girder3 of phase2 project) is at 7uA. Yesterday I did a leak check and all is well. Beamline leak free. Need to ohm out the harp and make sure wires are intact and not shorting out during motion. Once verified OK, we can add pinchoffs to all the RA valves of the phase2 beamline, and put the pump cart away. Assuming harp is good, we are done with the phase2 vacuum work upstream of the booster.

COMMENT ON 53-CH-1 Replace Condenser Fan Motor (INJ Chiller System)

Log Number: 4159060Logbook(s): Body: 

Spoke with owner about this task. The new LOTO protocol with be followed before work begins. During the execution of this task, cooling down in the INJ segment will not be interrupted. This work is planned to start between 1200-1500.

A HI review will need to be conducted before then. Consider this work approved once the review is complete.

Pages